2. A salesperson contacts ten potential customers. From past experience, we know that for each potential
customer, the probability of them making a purchase is .10.
[1] a. What is the probability the salesperson will make exactly one sale?


[1] b. What is the probability of making no sales?



[1] c. What is the probability the salesperson will make at least one sale?



[1] d. What is the expected number of sales?


[1] e. What is the standard deviation for the number of sales?

Answer by Boreal(15235)

binomial with n=10 and p=0.1
expected value is that product or 1
variance is 1*(1-p)=.9
sd is sqrt of 0.9 or 0.949
probability of 1 sale is 10*0.1^1*0.9^9=0.3874
probability of no sales is 0.9^10=0.3487
probability of at least 1 sale is 1-prob of no sales or 0.6513